Niles Co. has the following data related to an item of inventory: Inventory, March 1 Purchase, March 7 Purchase, March 16 InvNiles Co. has the following data related to an item of inventory: Inventory, March 1 400 units @ $2.10 Purchase, March 7 1400 units @ $2.20 Purchase, March 16 280 units @ $2.50 Inventory, March 31 520 units The value assigned to cost of goods sold if Niles uses FIFO is
Answer #1

Units sold = Units available for sale - Units in ending inventory

= (400+1400+280) - 520

= 1560 units.

Under Fifo, first available units are sold first. So

Calculate 1560 units from beginning

= (400*2.10) + (1560 - 400)*2.20

= (400*2.10) + (1160*2.20)

= 840 + 2552

= 3392